OpenStudy (chaseffs):

Solve for c. c + 2/5 = -1/10 Answers: c = -1/5 c = -1/2 c = 1/10 c = 3/5

sam (.sam.):

\[c + 2/5 = -1/10 \\ \\ c + \frac{2}{5} \color{blue}{-\frac{2}{5}}= -\frac{1}{10} \color{blue}{-\frac{2}{5}} \\ \\ c=-\frac{1}{10} \color{blue}{-\frac{2}{5}}\]

sam (.sam.):

Then common denominator

OpenStudy (chaseffs):

Fractions are my downfall ;-;

OpenStudy (chaseffs):

hold on. I might be onto something. working it out...

OpenStudy (chaseffs):

what I got was three Fifths.

OpenStudy (dan815):

1/5 = 2/10 2/5 =??

OpenStudy (dan815):



OpenStudy (dan815):